# Near-Optimal Adaptive Transforms for LLM Quantization

### v1

@@ 3.1 Algorithm
3
@@
5for $i\leftarrow 1\ \mathrm{to}\ d_{\mathrm{in}}/d\ \mathrm{(in\ parallel)}$ do
6
7 $j_{1},j_{2}\leftarrow\left(i-1\right)d+1,id$
[…]
8
[…]
9
 $\bm{W}^{\prime}\leftarrow\operatorname{Cholesky}\left(\bm{M}_{\mathrm{W}}\right)$ // $\bm{W}^{\prime}\bm{W}^{\prime\top}=\bm{M}_{\mathrm{W}}$
10
 $\bm{U},\bm{\Lambda}\leftarrow\operatorname{Eigendecomposition}\left(\bm{W}^{\prime\top}\bm{M}_{\mathrm{X}}\bm{W}^{\prime}\right)$ // $\bm{U}\bm{\Lambda}\bm{U}^{\top}=\bm{W}^{\prime\top}\bm{M}_{\mathrm{X}}\bm{W}^{\prime}$
11
 ${\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}\leftarrow\bm{H}\bm{\Lambda}^{-\frac{1}{4}}\bm{U}^{\top}\bm{W}^{\prime\top}$ // WUSH transform
12
 $\hat{\bm{W}}_{\left(i\right)}=q\left({\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}^{-\top}\bm{W}[j_{1}:j_{2},1:d_{\mathrm{out}}]\right)$ // pre-quantize weight block
13
14 end for
15
Algorithm 1 Compute WUSH Transforms and Pre-Quantize Weights
Algorithm˜1 summarizes the procedure used to compute the blockwise transforms and pre-quantized weights for each linear layer. In practice, the required second-order information can be obtained either from the calibration activations or from the Hessian matrix; the latter coincides, up to a constant factor, with the Hessian used in GPTQ. As a result, the calibration pipeline closely follows that of GPTQ: layers are processed sequentially, and after processing a layer, the calibration activations are propagated through the quantized layer to provide the inputs for calibrating the next one.
During inference, the forward pass is calculated as
$\sum_{i=1}^{d_{\mathrm{in}}/d}\hat{\bm{W}}_{\left(i\right)}^{\top}q\left({\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}\bm{X}_{\left(i\right)}\right).$ (25)
with the symbol $\bm{X}$ being overloaded as the new activations instead of the calibration ones.
3.2 Layer Loss
Table 1: Weight-activation RTN quantization losses with identity (I), random rotation (R), Hadamard (H), WUSH without Hadamard (WUS), and WUSH transforms.
Format Transform Q K V O Gate Up Down
MXFP4 I .01105 .01201 .01068 .00435 .00710 .00656 .00547
R .00761 .00773 .00914 .00384 .00556 .00568 .00407
H .00724 .00720 .00860 .00379 .00545 .00561 .00390
WUS .00627 .00722 .00405 .00357 .00576 .00475 .00446
WUSH .00334 .00334 .00330 .00276 .00449 .00439 .00339
NVFP4 I .00423 .00435 .00437 .00234 .00349 .00341 .00241
R .00498 .00501 .00586 .00254 .00363 .00368 .00266
H .00560 .00562 .00670 .00258 .00371 .00379 .00278
WUS .00226 .00236 .00230 .00200 .00304 .00301 .00223
WUSH .00240 .00244 .00228 .00192 .00309 .00302 .00233
INT4 I .17004 .12277 .01315 .00455 .05585 .00983 .01934
R .00579 .00584 .00696 .00294 .00422 .00429 .00310
H .00557 .00555 .00680 .00286 .00409 .00425 .00303
WUS .21268 .14233 .01067 .00454 .05021 .00742 .01308
WUSH .00239 .00243 .00254 .00210 .00343 .00343 .00255
We report the layerwise weight-activation RTN quantization loss ($L_{2}$ loss normalized by the number of elements) for each linear layer in the 18th transformer block of Qwen3-8B, using 32 calibration samples of sequence length 2048 from the FineWeb-Edu (NEURIPS2024_370df50c) dataset. We compare the WUSH transform with the identity (I), random rotation (R) averaged over 10 runs, Hadamard (H), and WUSH without Hadamard (WUS). Table˜1 summarizes the losses for MXFP4, NVFP4, and INT4 formats, where INT4 denotes 4-bit integers with Gaussian MSE clipping and BF16 scales of group size 32. Across formats, WUSH substantially reduces the loss: WUSH consistently yields the smallest loss for MXFP4 and INT4, while WUSH and WUS are almost equally optimal for NVFP4.
These empirical trends are broadly consistent with our theoretical analysis, suggesting that the approximation assumptions (Section˜2.1) and the smooth error models (Sections˜2.2.1 and 2.3.1) capture the dominant sources of quantization loss in practice. At the same time, MXFP4 behaves as a hybrid between ideal FP and INT quantization: the effective mantissa step changes uniformly, imparting INT-like behavior, especially in the subnormal regime and small exponent ranges. This helps explain why Hadamard transforms still provide measurable gains for the MXFP4 format, despite the ideal FP theory (Section˜2.2.3) predicting no benefit for purely orthogonal transforms.
3.3 LLM Benchmarks
We now present preliminary experiments evaluating the accuracy of models using the WUSH transform, relative to the identity (standard RTN) and Hadamard. We conduct our experiment on the LLama-3.2-3B-Instruct, LLama3.1-8B-Instruct, Qwen3-8B, and Qwen3-14B models. We use Platinumbench (vendrow2025large) and the LM Eval Harness (gao2021framework) for evaluation.
The full results for LM Eval Harness tasks are shown in Table˜2. Here, we observe that WUSH achieves the highest average recovery score among all the methods tested, often by large margins. Remarkably, on large models such as Qwen3-14B, it helps narrow the accuracy gap between MXFP4 and NVFP4 to nearly 0.5%.
We plot the results for Platinumbench in Figures˜3, 4, 1 and 2. Again, we observe that WUSH provides significant accuracy improvements, especially in the case of MXFP.
Table 2: Accuracy results on the LM Eval Harness for round-to-nearest (RTN) quantization of weights and activations with identity (I), Hadamard (H), and WUSH transforms.
Model Format Transform MMLU GSM8K HellaSwag WinoGrande Avg. Recovery%
Llama-3.2-3B-Instruct BF16 - 64.43 78.01 73.42 70.09 71.49
NVFP4 I 61.02 71.04 70.90 66.77 67.43 94.33
H 59.91 64.82 69.77 65.59 65.02 90.95
WUSH 61.25 71.78 71.08 67.18 67.82 94.87
MXFP4 I 56.81 60.80 67.30 64.56 62.37 87.25
H 55.99 61.11 68.36 64.09 62.39 87.27
WUSH 58.45 68.60 69.99 66.54 65.89 92.17
Llama-3.1-8B-Instruct BF16 - 72.76 85.06 80.01 77.90 78.93
NVFP4 I 68.64 80.29 77.97 74.43 75.33 95.44
H 66.96 77.41 77.29 74.35 74.00 93.75
WUSH 68.83 78.57 78.22 75.47 75.28 95.37
MXFP4 I 62.21 69.52 73.76 72.61 69.53 88.08
H 63.03 75.44 75.51 71.90 71.47 90.54
WUSH 66.85 75.16 77.28 73.56 73.21 92.75
Qwen3-8B BF16 - 72.98 90.90 75.52 70.56 77.49
NVFP4 I 70.47 88.40 74.44 69.53 75.71 97.70
H 70.19 86.35 73.02 68.11 74.42 96.03
WUSH 70.86 89.33 74.82 69.79 76.20 98.33
MXFP4 I 67.69 84.23 71.24 67.40 72.64 93.74
H 67.53 87.04 71.48 67.64 73.42 94.75
WUSH 70.29 87.95 74.06 69.50 75.45 97.36
Qwen3-14B BF16 - 77.18 91.96 79.84 74.27 80.81
NVFP4 I 75.20 90.22 77.38 73.32 79.03 97.80
H 74.98 92.04 77.76 72.38 79.29 98.12
WUSH 75.43 91.78 78.82 73.18 79.80 98.75
MXFP4 I 72.92 90.22 76.68 71.51 77.83 96.31
H 73.13 87.41 76.29 71.67 77.13 95.44
WUSH 74.56 91.87 78.18 73.15 79.44 98.30
Figure 1: Comparison of different transforms on Qwen3-8B for both NVFP4 and MXFP4 RTN quantization on Platinum benchmark tasks. The left table shows accuracy results across the individual Platinum benchmark tasks, while the right plot shows the average accuracy scores together with their standard deviations for each transform.
Figure 2: Comparison of different transforms on Qwen3-14B for both NVFP4 and MXFP4 RTN quantization on Platinum benchmark tasks. The left table shows accuracy results across the individual Platinum benchmark tasks, while the right plot shows the average accuracy scores together with their standard deviations for each transform.
Figure 3: Comparison of different transforms on Llama3.2-3B Instruct for both NVFP4 and MXFP4 RTN quantization on Platinum benchmark tasks. The left table shows accuracy results across the individual Platinum benchmark tasks, while the right plot shows the average accuracy scores together with their standard deviations for each transform.
Figure 4: Comparison of different transforms on Llama3.1-8B Instruct for both NVFP4 and MXFP4 RTN quantization on Platinum benchmark tasks. The left table shows accuracy results across the individual Platinum benchmark tasks, while the right plot shows the average accuracy scores together with their standard deviations for each transform.
4 Future Work
A natural next step is to integrate WUSH with advanced post-training quantization schemes such as GPTQ. In this setting, the weights are updated iteratively through error propagation, while WUSH is constructed from the second-order statistics of the updated weights. Designing an algorithm that jointly updates the weights and the associated blockwise transforms is non-trivial and will require adapting both the optimization schedule and the calibration pipeline.
From a systems perspective, efficient kernel support for WUSH is another important direction. Unlike blockwise Hadamard rotations, which reuse a single transform across all blocks, WUSH generally assigns a distinct transform to each block. This per-block specialization complicates kernel reuse and fusion in existing low-level libraries.
Finally, it will be useful to explore structurally simplified variants of WUSH that trade optimality for lower computational overhead. One promising avenue is to approximate the data-aware component with a diagonal matrix, thereby reducing the cost of applying the transforms at inference time.
5 Conclusion
We have studied blockwise weight-activation quantization for LLMs in the presence of heavy-tailed outliers and have shown that, under mild assumptions, this problem admits closed-form linear transforms for both FP and INT block quantizers. Our construction, WUSH, combines a fixed Hadamard backbone with a data-aware component derived from second-order statistics, yielding a non-orthogonal blockwise transform that is provably optimal and remains amenable to efficient implementation. These results clarify the empirical success of Hadamard rotations and indicate that principled, data-aware block transforms can significantly narrow the gap between low-bit and full-precision models.
Acknowledgment
Appendix A Maximum Inequalities for Tail-Bounded Distributions
@@
Let $X\geq 0$ be a random variable with the density function $f$ and the survival function $S(t)=\mathbb{P}\left(X\geq t\right)$ with $t\geq 0$. Let $g:\left[0,\infty\right)\to\mathbb{R}$ be integrable and define $G\left(t\right)=\int_{0}^{t}g\left(u\right)\mathrm{d}u$. Assume that $\lim_{t\to\infty}G\left(t\right)S\left(t\right)=0$. Then $\mathbb{E}G\left(X\right)=\int_{0}^{\infty}g\left(t\right)S\left(t\right)\mathrm{d}t$.
Proof Since $X$ has density $f$ and takes values in $0,\infty)$, $\mathbb{E}G\left(X\right)=\int_{0}^{\infty}G\left(t\right)f\left(t\right)\mathrm{d}t$. The survival function satisfies $S\left(t\right)=\mathbb{P}\left(X\geq t\right)=\int_{t}^{\infty}f(u)\mathrm{d}u$, so $S$ is differentiable almost everywhere with $S^{\prime}\left(t\right)=-f\left(t\right)$. Hence, $\mathbb{E}G\left(X\right)=\int_{0}^{\infty}G\left(t\right)f\left(t\right)\mathrm{d}t=-\int_{0}^{\infty}G\left(t\right)S^{\prime}\left(t\right)\mathrm{d}t$. We now integrate by parts, using $G^{\prime}\left(t\right)=g\left(t\right)$,
@@
By definition, $G\left(0\right)=0$, and by assumption $\lim_{t\to\infty}G\left(t\right)S\left(t\right)=0$, the boundary term $\left[G\left(t\right)S\left(t\right)\right]_{0}^{\infty}=0$. Therefore, $\mathbb{E}G\left(X\right)=\int_{0}^{\infty}g\left(t\right)S\left(t\right)\mathrm{d}t$.
A.1 Gaussian Distribution
@@
A probability distribution of a random variable $X\in\mathbb{R}$ is called sub-Gaussian with variance proxy $\sigma^{2}$ if $\mathbb{E}\exp\left(\left(X-\mathbb{E}X\right)t\right)\leq\exp\left(2^{-1}\sigma^{2}t^{2}\right)$ for all $t\in\mathbb{R}$, and the smallest such $\sigma^{2}$ is called the optimal variance proxy.
@@
Chernoff bound on the tail: $\mathbb{P}\left(\left|X-\mathbb{E}X\right|\geq t\right)\leq 2\exp\left(-2^{-1}\sigma^{-2}t^{2}\right)$ for all $t\geq 0$.
@@
Let $X_{1},\dots,X_{d}$ be zero-mean Gaussian random variables with variance $\sigma_{1}^{2},\dots,\sigma_{d}^{2}$. Denote $\sigma_{\star}^{2}=\max_{k}\sigma_{k}^{2}=\max_{k}\mathbb{E}X_{k}^{2}$ as the smallest variance proxy for $X_{1},\dots,X_{d}$. Using Lemma 5,
$\ \mathbb{E}\max_{k}X_{k}^{2}=2\int_{0}^{\infty}t\ \mathbb{P}\left(\max_{k}\left|X_{k}\right|\geq t\right)\mathrm{d}t=2\int_{0}^{\infty}t\ \mathbb{P}\left(\bigcup_{k=1}^{d}\left\{\left|X_{k}\right|\geq t\right\}\right)\mathrm{d}t$ (27)
@@
$\leq$ $2\int_{0}^{t_{\star}}t\ \mathrm{d}t+4d\int_{t_{\star}}^{\infty}t\exp\left(-2^{-1}\sigma_{\star}^{-2}t^{2}\right)\mathrm{d}t=\left[t^{2}\right]_{0}^{t_{\star}}+4d\left[-\sigma_{\star}^{2}\exp\left(-2^{-1}\sigma_{\star}^{-2}t^{2}\right)\right]_{t_{\star}}^{\infty}$
@@
without needing the independence of $X_{k}$. Also, considering the trivial bound $\mathbb{E}\max_{k}X_{k}^{2}\leq\mathbb{E}\sum_{k=1}^{d}X_{k}^{2}=\sum_{k=1}^{d}\mathbb{E}X_{k}^{2}\leq d\max_{k}\mathbb{E}X_{k}^{2}$, then $\mathbb{E}\max_{k}X_{k}^{2}\leq\min\left\{d,\left(2\ln\left(2d\right)+2\right)\right\}\max_{k}\mathbb{E}X_{k}^{2}$.
A.2 Laplacian Distribution

### v2

@@ A.3 Equation Derivations
For orthogonal $\bm{T}$, we have $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}=d$ and $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\mathbb{E}_{\bm{y}}\left\|\bm{y}\right\|^{2}$. Using Equation˜36, $\mathbb{E}_{\bm{y}}\left\|\bm{y}\right\|^{2}=\operatorname{tr}\left(\bm{S}^{2}\right)$. By Equation˜50,
$\ \operatorname{tr}\left(\bm{S}^{2}\right)=d^{-1}\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}$ (58)
$\leq$ $\ \left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\leq\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=d\operatorname{tr}\left(\bm{S}^{2}\right).$
A.4 Quantization Error Modeling Justifications
A.4.1 Integer (INT) Types
We model the quantization error using pseudo noise similar to DiffQ (Défossez et al., 2022).
Define the scalar quantizer of $b$-bit symmetric integer as
$q\left(x\right)=\left(\left\lfloor xs^{-1}\right\rfloor+2^{-1}\right)s,$ (59)
where $x\in\mathbb{R}$, scale $s\in\mathbb{R}_{\neq 0}$, and $xs^{-1}\in\left[-2^{b-1},2^{b-1}\right)$.
To simplify the analysis, we model the quantizer as
$q\left(x\right)=\left(xs^{-1}+\xi\right)s=x+s\xi$ (60)
with random noise $\xi\sim\mathrm{Uniform}\left(-2^{-1},2^{-1}\right)$. Then, the quantization error can be expressed as
$q\left(x\right)-x=s\xi.$ (61)
For a vector $\bm{\alpha}\in\mathbb{R}^{d}$, we use the AbsMax scale $s=2^{1-b}\left\|\bm{\alpha}\right\|_{\infty}$. Define
$\eta=2^{1-b}\xi,$ (62)
then we have $s\xi=\eta\left\|\bm{\alpha}\right\|_{\infty}$. The quantization error of vector $\bm{\alpha}$ is
$\varepsilon\left(\bm{\alpha}\right)=\left\|\bm{\alpha}\right\|_{\infty}\bm{\eta}$ (63)
with $\bm{\eta}=\left[\eta_{1},\dots,\eta_{d}\right]^{\top}\in\mathbb{R}^{d}$ being a random vector of i.i.d. samples $\eta\sim\mathcal{D}_{\mathrm{\eta}}$ and
$\mathbb{E}_{\eta}\eta=2^{1-b}\mathbb{E}_{\xi}\xi=0$ (64)
A.4.2 Floating-Point (FP) Types
The FP format $\mathrm{E}a\mathrm{M}b$ is defined as a tuple $\left(s,e,m\right)$ with a sign $s\in\left\{0,1\right\}$, an exponent $e\in\left\{0,\dots,2^{a}-1\right\}$, and a mantissa $m\in\left\{0,\dots,2^{b}-1\right\}$. Assume we do not need to consider the subnormal and special number cases. The tuple $\left(s,e,m\right)$ represents the number
$x_{\mathrm{E}a\mathrm{M}b}=\left(-1\right)^{s}2^{e-2^{a-1}+1}\left(1+2^{-b}m\right).$ (65)
Figure 5: The two curves $y=1+x$ and $y=2^{x}$ are close for $0\leq x\leq 1$.
Using the approximation $1+\left(\cdot\right)\approx 2^{\left(\cdot\right)}$ for $2^{-b}m\in\left[0,1\right]$ (visualized in Figure˜5), we define a smoothed version of $\mathrm{E}a\mathrm{M}b$, named $\mathrm{SE}a\mathrm{M}b$, as
$x_{\mathrm{SE}a\mathrm{M}b}=\left(-1\right)^{s}2^{e-2^{a-1}+1}2^{2^{-b}m}=\left(-1\right)^{s}2^{2^{-b}\left(2^{b}e+m\right)-2^{a-1}+1}$ (66)
with
$2^{b}e+m=2^{b}\left(\log_{2}\left|x_{\mathrm{SE}a\mathrm{M}b}\right|+2^{a-1}-1\right)$ (67)
being an $\left(a+b\right)$-bit unsigned integer (concatenating the binary representations of $e$ and $m$). The minimum and maximum of $\left|x_{\mathrm{SE}a\mathrm{M}b}\right|$ are
$\min\left|x_{\mathrm{SE}a\mathrm{M}b}\right|=$ $2^{-2^{a-1}+1},$ (68)
$\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|=$ $2^{2^{-b}\left(2^{a+b}-1\right)-2^{a-1}+1}=2^{2^{a-1}-2^{-b}+1},$
respectively.
Denote $\mathrm{SE}a\mathrm{M}b\left(x\right)$ as the quantizer that casts $x\in\left[-\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|,\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|\right]$ to the nearest number in the $\mathrm{SE}a\mathrm{M}b$ type. Assume $e\in\left(-\infty,2^{a}-1\right]$ so that the smaller values can be represented more precisely. $\mathrm{SE}a\mathrm{M}b\left(x\right)$ can be approximated using integer rounding in logarithmic space and first-order expansion. For $x\neq 0$,
$\mathrm{SE}a\mathrm{M}b\left(x\right)\approx$ $\ \mathrm{sgn}\left(x\right)2^{2^{-b}\left\lfloor 2^{b}\left(\log_{2}\left|x\right|+2^{a-1}-1\right)\right\rceil-2^{a-1}+1}$ (69)
$=$ $\ \mathrm{sgn}\left(x\right)2^{2^{-b}\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil}$
For $x=0$,
$\mathrm{SE}a\mathrm{M}b\left(0\right)\approx\lim_{x\to 0}x\left(1+\left(\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil-2^{b}\log_{2}\left|x\right|\right)2^{-b}\ln 2\right)=0.$ (70)
We model the casting error as
$\mathrm{SE}a\mathrm{M}b\left(x\right)-x=\left(\ln 2\right)2^{-b}\xi x$ (71)
where $x\in\left[-\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|,\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|\right]$ and, similar to the INT types in Equation˜60, the random variable $\xi\sim\mathrm{Uniform}\left(-2^{-1},2^{-1}\right)$.
@@ A.4.2 Floating-Point (FP) Types
$q\left(x\right)=$ $\ \mathrm{SE}a\mathrm{M}b\left(xs^{-1}\right)\mathrm{SE}\tilde{a}\mathrm{M}\tilde{b}\left(s\right)$ (72)
$=$ $\ \left(1+\left(\ln 2\right)2^{-b}\xi\right)xs^{-1}\left(1+\left(\ln 2\right)2^{-\tilde{b}}\tilde{\xi}\right)s$
Define
$\eta=\left(1+\left(\ln 2\right)2^{-b}\xi\right)\left(1+\left(\ln 2\right)2^{-\tilde{b}}\tilde{\xi}\right)-1.$ (73)
Then, the quantization error can be expressed as
$q\left(x\right)-x=\eta x.$ (74)
And the expected quantization error is 0 because
A.5 Maximum Inequalities for Tail-Bounded Distributions
@@
Let $X\geq 0$ be a random variable with the density function $f$ and the survival function
$S(t)=\mathbb{P}\left(X\geq t\right)$ (76)
with $t\geq 0$. Let $g:\left[0,\infty\right)\to\mathbb{R}$ be integrable and define
$G\left(t\right)=\int_{0}^{t}g\left(u\right)\mathrm{d}u.$ (77)
Assume that
$\lim_{t\to\infty}G\left(t\right)S\left(t\right)=0.$ (78)
Then,
$\mathbb{E}G\left(X\right)=\int_{0}^{\infty}g\left(t\right)S\left(t\right)\mathrm{d}t.$ (79)
Proof Since $X$ has density $f$ and takes values in $0,\infty)$,
$\mathbb{E}G\left(X\right)=\int_{0}^{\infty}G\left(t\right)f\left(t\right)\mathrm{d}t.$ (80)
The survival function satisfies
$\begin{aligned} \end{aligned}S\left(t\right)=\mathbb{P}\left(X\geq t\right)=\int_{t}^{\infty}f\left(u\right)\mathrm{d}u,$ (81)
so $S$ is differentiable almost everywhere with $S^{\prime}\left(t\right)=-f\left(t\right)$. Hence,
$\mathbb{E}G\left(X\right)=\int_{0}^{\infty}G\left(t\right)f\left(t\right)\mathrm{d}t=-\int_{0}^{\infty}G\left(t\right)S^{\prime}\left(t\right)\mathrm{d}t.$ (82)
We now integrate by parts, using $G^{\prime}\left(t\right)=g\left(t\right)$,
@@
By definition, $G\left(0\right)=0$, and by assumption $\lim_{t\to\infty}G\left(t\right)S\left(t\right)=0$, the boundary term $\left[G\left(t\right)S\left(t\right)\right]_{0}^{\infty}=0$. Therefore,
$\mathbb{E}G\left(X\right)=\int_{0}^{\infty}g\left(t\right)S\left(t\right)\mathrm{d}t.$ (84)
A.5.1 Gaussian Distribution
@@
A probability distribution of a random variable $X\in\mathbb{R}$ is called sub-Gaussian with variance proxy $\sigma^{2}$ if
$\mathbb{E}\exp\left(\left(X-\mathbb{E}X\right)t\right)\leq\exp\left(2^{-1}\sigma^{2}t^{2}\right)$ (85)
for all $t\in\mathbb{R}$, and the smallest such $\sigma^{2}$ is called the optimal variance proxy.
@@
Chernoff bound on the tail:
$\mathbb{P}\left(\left|X-\mathbb{E}X\right|\geq t\right)\leq 2\exp\left(-2^{-1}\sigma^{-2}t^{2}\right)$ (86)
for all $t\geq 0$.
@@
Let $X_{1},\dots,X_{d}$ be zero-mean Gaussian random variables with variance $\sigma_{1}^{2},\dots,\sigma_{d}^{2}$. Denote
$\sigma_{\star}^{2}=\max_{k}\sigma_{k}^{2}=\max_{k}\mathbb{E}X_{k}^{2}$ (87)
as the smallest variance proxy for $X_{1},\dots,X_{d}$. Using Lemma 4,
$\mathbb{E}\max_{k}X_{k}^{2}=$ $2\int_{0}^{\infty}t\ \mathbb{P}\left(\max_{k}\left|X_{k}\right|\geq t\right)\mathrm{d}t$ (88)
$=$ $2\int_{0}^{\infty}t\ \mathbb{P}\left(\bigcup_{k=1}^{d}\left\{\left|X_{k}\right|\geq t\right\}\right)\mathrm{d}t$
@@
$\leq$ $2\int_{0}^{t_{\star}}t\ \mathrm{d}t+4d\int_{t_{\star}}^{\infty}t\exp\left(-2^{-1}\sigma_{\star}^{-2}t^{2}\right)\mathrm{d}t$
$=$ $\ \left[t^{2}\right]_{0}^{t_{\star}}+4d\left[-\sigma_{\star}^{2}\exp\left(-2^{-1}\sigma_{\star}^{-2}t^{2}\right)\right]_{t_{\star}}^{\infty}$
@@
without needing the independence of $X_{k}$. Also, considering the trivial bound
$\mathbb{E}\max_{k}X_{k}^{2}\leq\mathbb{E}\sum_{k=1}^{d}X_{k}^{2}=\sum_{k=1}^{d}\mathbb{E}X_{k}^{2}\leq d\max_{k}\mathbb{E}X_{k}^{2},$ (90)
then,
A.5.2 Laplacian Distribution